OpenStudy (anonymous):

A projectile is thrown upwards so that its distance above the ground after (t) seconds is h(t) = -16t^2+320t. After how many seconds does it reach its maximum height.

OpenStudy (anonymous):

-16t^2+320t=0 -16t(t-20)=0 t=20

OpenStudy (anonymous):

8, 10, 24, 32, 56, 78, 96, 107 seconds are all possible answers

OpenStudy (amistre64):

-32t +320 = 0 at t = 320/32 = 10